what is the slope-intercept form of a line that passes through points (2, 11) and (4, 17)?
Answer by lwsshak3(11628)

what is the slope-intercept form of a line that passes through points (2, 11) and (4, 17)?
***
standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
slope,m=∆y/∆x=(17-11)/(4-2)=6/2=3
equation: y=3x+b
solve for b using coordinates of one of the given points (2,11)
11=3*2+b
b=5
equation: y=3x+5